The solution set of the equation 4log9 x  -6.xlog9 2 + 8 = 0 is ______.

Grade:12

2 Answers

Y RAJYALAKSHMI
9 years ago
We have xlogay = ylogax
S, xlog9= 2log9x
4log9 x  -6.xlog2 + 8 = 0
=> (22)log9x – 6. 2log9x + 8 = 0
=> (2log9)2  - 6. 2log9x + 8 = 0
let 2log9x = y
Then, the given equation becomes, y2 – 6y + 8 = 0 and the solution is y = 4 & 2
when y = 4,  2log9x = 22  => log9x = 2 => x = 9= 81
when y = 2,  2log9x = 2  => log9x = 1 => x = 9
 
Ans:  Solution set = {9, 81}
 
 
 
 
 
Mostafijur Rahaman
9 years ago
thnx